# Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization

### v5

@@ A.3. Complexity Analysis on Coverage Axis
Figure A8. Distribution of randomly sampled candidate skeletal points in two different shapes. $(a)$ tubular structure. $(b)$ planar structure.
Table A5. Quantitative comparison on run time and shape approximation error between Coverage Axis and Coverage Axis++ when the number of surface samples is set as $5000$. The inputs are meshes. Time: Runtime measured in seconds. $|V|$: The number of skeletal points. $\overrightarrow{\epsilon}$: One-sided HD from surface to reconstruction. $\overleftarrow{\epsilon}$: One-sided HD from reconstruction to surface. $\overleftrightarrow{\epsilon}$: Two-sided HD between original surface and reconstruction.
@@
Bear-1
70
115.2 $3.651\%$ $4.958\%$ $4.958\%$ $5.1$ $2.752\%$ $2.795\%$ $2.795\%$
Bear-2
38
7.5 $7.412\%$ $7.385\%$ $7.412\%$ $2.5$ $4.552\%$ $5.997\%$ $5.997\%$
Bird
101
¿1000 $1.959\%$ $1.857\%$ $1.959\%$ $9.8$ $2.177\%$ $2.438\%$ $2.438\%$
Bug
145
¿1000 $5.701\%$ $5.636\%$ $5.701\%$ $12.2$ $3.074\%$ $3.005\%$ $3.074\%$
Camel
116
25.6 $2.488\%$ $2.238\%$ $2.488\%$ $9.5$ $2.513\%$ $2.936\%$ $2.936\%$
Chair
137
¿1000 $1.920\%$ $1.871\%$ $1.920\%$ $13.6$ $1.909\%$ $2.105\%$ $2.105\%$
Crab-1
92
43.9 $3.738\%$ $3.293\%$ $3.738\%$ $9.5$ $2.577\%$ $2.738\%$ $2.738\%$
Crab-2
83
210.5 $7.705\%$ $6.223\%$ $7.705\%$ $8.9$ $3.168\%$ $3.212\%$ $3.212\%$
Cup
286
¿1000 $9.176\%$ $5.026\%$ $9.176\%$ $27.8$ $9.174\%$ $6.390\%$ $9.174\%$
Dog
65
7.6 $4.557\%$ $4.386\%$ $4.557\%$ $5.0$ $2.642\%$ $2.876\%$ $2.876\%$
Elephant
89
147.1 $6.015\%$ $7.221\%$ $7.221\%$ $8.5$ $3.168\%$ $3.925\%$ $3.925\%$
Fandisk
210
¿1000 $2.925\%$ $2.778\%$ $2.925\%$ $23.7$ $2.613\%$ $3.057\%$ $3.057\%$
Femur
27
13.6 $2.842\%$ $2.782\%$ $2.842\%$ $2.0$ $2.035\%$ $2.035\%$ $2.035\%$
Fish
72
¿1000 $1.769\%$ $2.271\%$ $2.271\%$ $5.7$ $1.994\%$ $2.722\%$ $2.722\%$
Giraffe
50
25.9 $5.841\%$ $5.853\%$ $5.853\%$ $5.1$ $2.627\%$ $5.267\%$ $5.267\%$
Guitar
63
¿1000 $2.040\%$ $4.312\%$ $4.312\%$ $7.4$ $2.232\%$ $2.383\%$ $2.383\%$
Hand
46
131.4 $2.874\%$ $2.607\%$ $2.874\%$ $4.2$ $3.315\%$ $2.884\%$ $3.315\%$
Horse
56
200.2 $5.174\%$ $6.042\%$ $6.042\%$ $4.6$ $3.608\%$ $5.246\%$ $5.246\%$
Lifebuoy
38
96.7 $4.667\%$ $4.196\%$ $4.667\%$ $2.8$ $2.52\%$ $3.307\%$ $3.307\%$
Neptune
120
206.8 $2.115\%$ $5.751\%$ $5.751\%$ $11.4$ $3.128\%$ $5.542\%$ $5.542\%$
Pig
51
16.6 $8.535\%$ $8.575\%$ $8.575\%$ $3.5$ $6.038\%$ $8.327\%$ $8.327\%$
Plane
84
¿1000 $1.708\%$ $2.275\%$ $2.275\%$ $8.7$ $1.965\%$ $2.533\%$ $2.533\%$
Pliers
32
13.5 $2.072\%$ $1.925\%$ $2.072\%$ $2.8$ $2.081\%$ $2.055\%$ $2.081\%$
Pot
152
¿1000 $3.570\%$ $3.100\%$ $3.570\%$ $14.4$ $3.504\%$ $3.690\%$ $3.690\%$
Rapter
74
¿1000 $3.322\%$ $6.545\%$ $6.545\%$ $6.2$ $3.322\%$ $6.634\%$ $6.634\%$
Rocker
130
¿1000 $3.632\%$ $3.496\%$ $3.632\%$ $14.4$ $3.777\%$ $2.790\%$ $3.777\%$
Seahorse
53
¿1000 $2.763\%$ $11.156\%$ $11.156\%$ $4.8$ $2.288\%$ $5.515\%$ $5.515\%$
Spectacles
112
¿1000 $1.982\%$ $1.952\%$ $1.982\%$ $14.4$ $1.660\%$ $1.653\%$ $1.660\%$
Vase
109
226.1 $5.260\%$ $6.587\%$ $6.587\%$ $9.2$ $3.132\%$ $3.192\%$ $3.192\%$
Wine-glass
335
¿1000 $22.680\%$ $7.904\%$ $22.680\%$ $34.6$ $21.310\%$ $8.549\%$ $21.310\%$
Average
516.3 4.670% 4.673% 5.448% $9.7$ 3.695% 3.860% $4.429\%$
Next, we provide more analysis of the complexity of the Coverage Axis. From Tab. 1 of the main paper, we observe that the running times of the Coverage Axis for certain models (such as Bird, Chair, Guitar, and Spectacles) are extremely long, indicating that solving SCP (Eq. 2) for these models is particularly challenging. Note that a common characteristic of these shapes is that they all have planar shapes. We illustrate why solving SCP for models with planar shapes is difficult.
Given an arbitrary shape, under the meaning of medial axis transform, it can be divided into two types of shapes: tubular and planar structures, as depicted in Fig. A8. Without loss of generality, we assume that the size and sample density of the two shapes are the same. In such cases, if the tubular shape contains $q$ points, the planar shape would have $q^{2}$ points. In the SCP formulation, the 00-$1$ decision vector $\mathbf{v}$ in Eq. 2 of the main paper is defined in a way that each entry indicates whether the corresponding candidate point is selected. Consequently, for the tubular shape, there are $2^{q}$ possible combinations, while for the planar shape, there are $2^{q^{2}}$ possible combinations. As a result, solving the SCP becomes significantly more challenging when the model involves planar shapes due to the exponential increase in the number of possible combinations. In comparison, Coverage Axis++ sidesteps solving an optimization problem with high computation complexity by designing a simple yet effective heuristic algorithm, taking coverage and uniformity into consideration, which thus achieves significant acceleration.
A.4. Limitations and Future Works
Coverage Axis++ approximates the whole shape by an abstraction of local geometry with the union of the local geometries forming the entire shape. One limitation of Coverage Axis++ is that its whole process is conducted without a hard constraint forcing the consistency between the selected medial balls and the shape surface (See coverage rate statistics in Sec. A.2), like that in Coverage Axis (Dou et al., 2022). Although local approximation has been evidenced by some previous works (Lin et al., 2021; Ge et al., 2023; Amenta et al., 2001), we acknowledge that the inclusion of a global constraint imposes a much stricter guarantee for the shape correspondence.
Another limitation is that as we keep the whole pipeline in a neat format, the uniformity of the skeletonization achieved by the Coverage Axis++ can occasionally be slightly inferior compared to that of the Coverage Axis. In future work, the uniformity scoring term in Algorithm 1 of the main paper can be further improved.

### v7

@@ Appendix D Complexity Analysis on Coverage Axis
Figure D1: Distribution of randomly sampled candidate skeletal points in two different shapes. $(a)$ tubular structure. $(b)$ planar structure.
Table D4: Quantitative comparison on run time and shape approximation error between Coverage Axis and Coverage Axis++ when the number of surface samples is set as $5000$. The inputs are meshes. Time: Runtime measured in seconds. $|V|$: The number of skeletal points. Note we set the number of selected skeletal points of our method the same as the Coverage Axis [DLX22] for a fair comparison.
@@
Bear-1 70 115.2 $3.651\%$ $4.958\%$ $4.958\%$ $5.1$ $2.752\%$ $2.795\%$ $2.795\%$
Bear-2 38 7.5 $7.412\%$ $7.385\%$ $7.412\%$ $2.5$ $4.552\%$ $5.997\%$ $5.997\%$
Bird 101 >1000 $1.959\%$ $1.857\%$ $1.959\%$ $9.8$ $2.177\%$ $2.438\%$ $2.438\%$
Bug 145 >1000 $5.701\%$ $5.636\%$ $5.701\%$ $12.2$ $3.074\%$ $3.005\%$ $3.074\%$
Camel 116 25.6 $2.488\%$ $2.238\%$ $2.488\%$ $9.5$ $2.513\%$ $2.936\%$ $2.936\%$
Chair 137 >1000 $1.920\%$ $1.871\%$ $1.920\%$ $13.6$ $1.909\%$ $2.105\%$ $2.105\%$
Crab-1 92 43.9 $3.738\%$ $3.293\%$ $3.738\%$ $9.5$ $2.577\%$ $2.738\%$ $2.738\%$
Crab-2 83 210.5 $7.705\%$ $6.223\%$ $7.705\%$ $8.9$ $3.168\%$ $3.212\%$ $3.212\%$
Cup 286 >1000 $9.176\%$ $5.026\%$ $9.176\%$ $27.8$ $9.174\%$ $6.390\%$ $9.174\%$
Dog 65 7.6 $4.557\%$ $4.386\%$ $4.557\%$ $5.0$ $2.642\%$ $2.876\%$ $2.876\%$
Elephant 89 147.1 $6.015\%$ $7.221\%$ $7.221\%$ $8.5$ $3.168\%$ $3.925\%$ $3.925\%$
Fandisk 210 >1000 $2.925\%$ $2.778\%$ $2.925\%$ $23.7$ $2.613\%$ $3.057\%$ $3.057\%$
Femur 27 13.6 $2.842\%$ $2.782\%$ $2.842\%$ $2.0$ $2.035\%$ $2.035\%$ $2.035\%$
Fish 72 >1000 $1.769\%$ $2.271\%$ $2.271\%$ $5.7$ $1.994\%$ $2.722\%$ $2.722\%$
Giraffe 50 25.9 $5.841\%$ $5.853\%$ $5.853\%$ $5.1$ $2.627\%$ $5.267\%$ $5.267\%$
Guitar 63 >1000 $2.040\%$ $4.312\%$ $4.312\%$ $7.4$ $2.232\%$ $2.383\%$ $2.383\%$
Hand 46 131.4 $2.874\%$ $2.607\%$ $2.874\%$ $4.2$ $3.315\%$ $2.884\%$ $3.315\%$
Horse 56 200.2 $5.174\%$ $6.042\%$ $6.042\%$ $4.6$ $3.608\%$ $5.246\%$ $5.246\%$
Lifebuoy 38 96.7 $4.667\%$ $4.196\%$ $4.667\%$ $2.8$ $2.52\%$ $3.307\%$ $3.307\%$
Neptune 120 206.8 $2.115\%$ $5.751\%$ $5.751\%$ $11.4$ $3.128\%$ $5.542\%$ $5.542\%$
Pig 51 16.6 $8.535\%$ $8.575\%$ $8.575\%$ $3.5$ $6.038\%$ $8.327\%$ $8.327\%$
Plane 84 >1000 $1.708\%$ $2.275\%$ $2.275\%$ $8.7$ $1.965\%$ $2.533\%$ $2.533\%$
Pliers 32 13.5 $2.072\%$ $1.925\%$ $2.072\%$ $2.8$ $2.081\%$ $2.055\%$ $2.081\%$
Pot 152 >1000 $3.570\%$ $3.100\%$ $3.570\%$ $14.4$ $3.504\%$ $3.690\%$ $3.690\%$
Rapter 74 >1000 $3.322\%$ $6.545\%$ $6.545\%$ $6.2$ $3.322\%$ $6.634\%$ $6.634\%$
Rocker 130 >1000 $3.632\%$ $3.496\%$ $3.632\%$ $14.4$ $3.777\%$ $2.790\%$ $3.777\%$
Seahorse 53 >1000 $2.763\%$ $11.156\%$ $11.156\%$ $4.8$ $2.288\%$ $5.515\%$ $5.515\%$
Spectacles 112 >1000 $1.982\%$ $1.952\%$ $1.982\%$ $14.4$ $1.660\%$ $1.653\%$ $1.660\%$
Vase 109 226.1 $5.260\%$ $6.587\%$ $6.587\%$ $9.2$ $3.132\%$ $3.192\%$ $3.192\%$
Wine-glass 335 >1000 $22.680\%$ $7.904\%$ $22.680\%$ $34.6$ $21.310\%$ $8.549\%$ $21.310\%$
Average - 516.3 4.670% 4.673% 5.448% $9.7$ 3.695% 3.860% $4.429\%$
Next, we provide more analysis of the complexity of the Coverage Axis. From Tab. 1 of the main paper, we observe that the running times of the Coverage Axis for certain models (such as Bird, Chair, Guitar, and Spectacles) are extremely long, indicating that solving SCP for these models is particularly challenging. Note that a common characteristic of these shapes is that they all have planar shapes. We illustrate why solving SCP for models with planar shapes is difficult.
Given an arbitrary shape, under the meaning of medial axis transform, it can be divided into two types of shapes: tubular and planar structures, as depicted in Fig. D1. Without loss of generality, we assume that the size and sample density of the two shapes are the same. In such cases, if the tubular shape contains $q$ points, the planar shape would have $q^{2}$ points. Recall that the SCP amounts to the following $0-1$ integer programming:
$\begin{split}&\min\left\|\mathbf{v}\right\|_{2},\\ &s.t.\ \begin{array}[]{c}\mathbf{Dv}\geq\mathds{1},\\ \end{array}\end{split}$ (10)
where $\mathds{1}$ is the vector of all ones, $\geq$ is applied element-wise, and $\mathbf{v}\in\{0,1\}^{n\times 1}$ is a decision vector, with its $i$-th element $v_{i}\in\{0,1\}$ indicating if the candidate skeletal point $p_{i}$ is selected. Consequently, for the tubular shape, there are $2^{q}$ possible combinations of $v$, while for the planar shape, there are $2^{q^{2}}$ possible combinations. As a result, solving the SCP becomes significantly more challenging when the model involves planar shapes due to the exponential increase in the number of possible combinations. In comparison, Coverage Axis++ sidesteps solving an optimization problem with high computation complexity by designing a simple yet effective heuristic algorithm, taking coverage and uniformity into consideration, which thus achieves significant acceleration.
Appendix E Skeletal Point Connectivity
In this Section, we discuss the necessity of mesh connectivity in the medial axis for various downstream applications. 1) Reconstruction: The connectivity of meshes and edges in the Medial Axis Transform (MAT) is important for shape reconstruction. This is because the connectivity allows for interpolation between medial axis balls, a.k.a., slab meshes, thereby achieving relatively higher accuracy in shape reconstruction. In this process, the connectivity between medial axis points is key to establishing this interpolation. If we disregard this connectivity, the results will exhibit obvious errors; See Fig. E2. 2) Shape Segmentation: As MAT contains a medial surface, previous unsupervised structural decomposition/shape segmentation [LLL20, LLL21] has shown that detecting dimensional changes and non-manifold branches on the medial mesh can be effective for the shape segmentation, e.g., use the information from manifold branch and dimensional change for segmentation; See Fig. E3 (a). 3) Hexmeshing: Based on the shape description ability of medial mesh, one could produce all-hex meshing using topological and geometrical information of the medial mesh, which is more general than previous methods based on curve-skeleton as pointed out by [ZXW24]; See Fig. E3 (b).
[…]
Figure E2: The medial mesh part among the skeletal points in the MAT enables interpolation between medial axis balls, resulting in higher accuracy shape reconstruction (e). In contrast, reconstruction using only skeletal points (c) or without a medial mesh connection (d) typically leads to larger approximation errors.
[…]
Figure E3: (a) The medial axis transform (MAT) includes a medial surface. Unsupervised shape segmentation methods show that detecting dimensional changes and non-manifold branches on the medial mesh effectively aids segmentation  [LLL20]. (b) In the application of all-hex meshing [ZXW24], the medial axis helps with computing the cross-field to guide the generation of the hexahedral mesh during the optimization.